OpenStudy (anonymous):

A(-1,3), B(2,5). Draw vectors AB and BA and give vectors in component form equivalent to each of these vectors.

OpenStudy (anonymous):

AB= (2-(-1)i + (5-3)j=3i+2j BA=(-1-2)i + (3-5)J = -3i-2j Draw both of these points on a graph and connect them.